The focal length of a lens of camera is 5f and that of another is 2.5f what is the time of exposure for the second if for the first one is 1/200 sec? a)1/200 b)1/800 c)1/6400 d)1/3200

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
f = focal length/aperature
 
f number decrease by 2
 
Time of exposure should decrese by 2² = 4
 
Hence time of exposure = (1/4)*(1/200) = 1/800
 
Regards
Arun (askIITians forum expert)